Can someone assist me in this problem? Thanks in advance.

A right triangle is a triangle with one angle measuring 90°. In a right triangle, the lengths of the sides are related by the Pythagorean Theorem,c2=a2+b2 , where c is the length of the hypotenuse (the side opposite the 90° angle). Find the length of the hypotenuse when the other two sides’ measurements are 6 feet and 8 feet

Found 2 solutions by checkley75, jayanthihemkos@yahoo.com:
Answer by checkley75(3666)

c^2=6^2+8^2
c^2=36+64
c^2=100
c=sqrt100
c=10 answer for the hypotenuse.

Answer by jayanthihemkos@yahoo.com(87)

Length of side a= 6 feet
Length of side b = 8 feet
Length of the hypotenuse c=?
According to pythagorean theorem, we have c2 = a2 + b2
plugging the values, we have
c2 = (6)square+(8)square
c2 = 36 + 64
c2= 100
c = square root of 100
c = 10 feet
hence the length of the hypotenuse = 10 feet

Jayanthi